Child pages
  • 201707 - Public report
Skip to end of metadata
Go to start of metadata

You are viewing an old version of this page. View the current version.

Compare with Current View Page History

Version 1 Next »

Unscheduled:

  • July 29th: Site SFA24, Installation-ID 100130048 on Assistentgatan 23, Skellefteå was down due to a power outage. The site owner has restored the power and the connection is up again.
  • July 18th: The link between Abisko and Rensjön was down due to a fiber break. The fiber has been repaired and connectivity is fully restored.
  • July 14th: BGP between our router in Filipstad and Stockholm was down due to provider emergency maintenance.
  • July 7th: Difficulties reaching connect.sunet.se due to network problems.

Scheduled:

  • July 31st: SUNET performed software upgrade on one of the core routers in Stockholm.
  • July 31st: The network interfaces on the transit router in Borås were restarted, one a the time, in order to correct a critical software bug.
  • July 30th: Carrier performed fiber maintenance. This affecedt the link between Arvidsjaur and Kåbdalis.
  • July 28th: SUNET performed upgrade on core routers.
  • July 27th: SUNET performed software upgrade on core routers. The routers are sbo-r1, lba-r1, trh-r1, hsd-r1, vxj-r1 and bos-r1
  • July 27th: Fiber supplier performed fiber optimization maintenance between Gällivare (GLV1) and Jokkmokk (JKK1).
  • July 17th: The fiber between the router located at Högskolan i Halmstad and Halmstad POP increased the attenuation value with more than 4 dbm. Fiber provider is performed fault trace the issue during above window but did not find any fault. Investigation will follow at the original trouble ticket.
  • July 12th: An emergency software maintenance was performed on the Sunet Survey service.
  • July 10th: SUNET performed software upgrade on router onsala-r1.sunet.se
  • July 10th: SUNET performed software upgrade on routers klx1-r1, hj-r1, hh-r2, ume8-r1 and uu-r1.
  • July 6th: A planned maintenance to migrate SUNET Adobe connect to a new setup of servers was performed.

Scope

Unscheduled

Scheduled

Total

Hardware

0 pcs

1 pcs 00:23

1 pcs 00:23

Link

1 pcs 00:08

2 pcs 12:45

3 pcs 12:53

None

3 pcs 23:48

3 pcs 08:55

6 pcs 1d 08:43

Routing

0 pcs

0 pcs

0 pcs

Software

0 pcs

5 pcs 03:38

5 pcs 03:38

  • No labels